the area of a square is calculated as (100+-2)cm2.the measured value of its length

Grade:11

4 Answers

Khimraj
5 years ago
error in area = (2/100)*100 = 2%
error in area = 2*(error in length of side)
so error in length of side = 1%
so length odf side is 10 \pm 0.1 cm
Hope it clears....................
Arun
5 years ago
Dear student
 
If area is 100 cm²
 
Then length = sqrt (100) = 10 cm
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Arun
5 years ago
Dear student
 
I forgot about error there will be an error of 1 % also as area has 2% error.
Hence
Length = 10+- 0.1 cm
Umesh lahoti
5 years ago
Maximum error in length=0.1cm
By error analysis-  2/100=2×0.1/x
Now x= 10
 
Measured value of length=(10+-0.1)